# VPGS-SLAM: Voxel-based Progressive 3D Gaussian SLAM in Large-Scale Scenes

### v1

@@ Abstract
3D Gaussian Splatting has recently shown promising results in dense visual SLAM. However, existing 3DGS-based SLAM methods are all constrained to small-room scenarios and struggle with memory explosion in large-scale scenes and long sequences. To this end, we propose VPGS-SLAM, the first 3DGS-based large-scale RGBD SLAM framework for both indoor and outdoor scenarios. We design a novel voxel-based progressive 3D Gaussian mapping method with multiple submaps for compact and accurate scene representation in large-scale and long-sequence scenes. This allows us to scale up to arbitrary scenes and improves robustness (even under pose drifts). In addition, we propose a 2D-3D fusion camera tracking method to achieve robust and accurate camera tracking in both indoor and outdoor large-scale scenes. Furthermore, we design a 2D-3D Gaussian loop closure method to eliminate pose drift. We further propose a submap fusion method with online distillation to achieve global consistency in large-scale scenes when detecting a loop. Experiments on various indoor and outdoor datasets demonstrate the superiority and generalizability of the proposed framework. The code will be open source on https://github.com/dtc111111/vpgs-slam.
@@
Figure 1: We present VPGS-SLAM, the first large-scale SLAM framework with voxel-based progressive 3D Gaussian representation, 2D-3D assisted camera tracking and 3D Gaussian loop closure. Depicted in the middle, we demonstrate the large-scale globally consistent 3D Gaussian map built with our approach. At the top and bottom of the figure, we include zoomed-in views of the map with RGB and depth images rendered by our method, indicated by dashed blue and yellow boxes.
1 Introduction
@@
Visual Simultaneous localization and mapping (SLAM) has been a fundamental computer vision problem with wide applications such as autonomous driving, robotics, and remote sensing. Several traditional methods orbslam2 ; deng ; xie ; cte have been introduced over the years. They use handcraft descriptors for image matching and represent scenes using sparse feature point maps. Due to the sparse nature of such point cloud, it is difficult for humans to understand how machines interact with the scene, and these methods cannot meet the demands of collision avoidance and motion planning. Attention then turns to dense scene reconstruction, exemplified by DTAM dtam , Kintinuous Kintinuous , and ElasticFusion elasticfusion . However, their performance is limited by high memory consumption, slow processing speeds.
@@
Following the introduction of Neural Radiance Fields (NeRF), numerous research efforts have focused on combining implicit scene representation with SLAM systems. iMAP imap pioneered the use of a single MLP to represent the scene, while niceslam ; eslam ; coslam ; plgslam have further enhanced scene representation through hybrid feature grids, axis-aligned feature planes, joint coordinate-parametric encoding, and multiple implicit submaps. To further improve rendering speed, recent methods have started to explore 3D Gaussian Splatting (3DGS) 3dgs in SLAM systems, as demonstrated by splatam ; monogs ; gaussianslam ; compactslam . GS-based SLAM methods leverage a point-based representation associated with 3D Gaussian attributes and adopt the rasterization pipeline to render the images, achieving fast rendering speed and promising image quality.
@@
However, existing SLAM systems primarily focus on small-scale indoor environments; they face significant challenges in representing large-scale scenes (e.g., multi-room apartments and urban scenes). We outline the key challenges for indoor and outdoor large-scale GS-based SLAM systems: a) Redundant 3D Gaussian ellipsoids: existing methods employ a substantial number of 3D Gaussian ellipsoids to represent the scene. These methods typically require more than 500MB to represent a small room-scale scene, which severely limits their applicability and results in memory overload in large-scale environments. b) Accumulation of errors and pose drift: Existing works rely on rendered loss for camera tracking, which proves inaccurate and unreliable in large-scale real-world environments with dramatic movement, motion blur and exposure change. c) Global inconsistency: In long sequences and large-scale scenes, when the robot revisits the same location, it is essential to ensure spatial correlations, and long-term memory for global consistency.
@@
To this end, we propose VPGS-SLAM, the first 3DGS-based large-scale SLAM framework with voxel-based progressive 3D Gaussian representation, 2D-3D fusion camera tracking, and 2D-3D Gaussian based loop detection and correction, and submap fusion. We use a collection of submaps to represent the entire scene, which dynamically initializes local scene representation when the camera moves to the bounds of the local submaps. The entire scene is divided into multiple local submaps, which can significantly improve the scene representation capacity of large-scale scenes. The submap parameters do not need to be retained in memory after optimization, which can significantly reduce online memory requirements and enhances the scalability of the framework.
@@
For camera tracking, we propose a 2D-3D fusion tracking method to combine the 3D Gaussian ellipsoids geometric information with 2D photometric information. We optimize the pose with 2D photometric loss using RGB and depth loss in the coarse-level optimization stage. We then incorporate a 3D voxel based ICP method to perform frame-to-map pose estimation in the fine-level stage. We provide a good initial estimate for 3D point matching through the 2D rendering loss, enabling further accurate pose estimation. To handle both indoor and outdoor environments, we propose an adaptive 2D information assessing and parameter selection strategy. By assessing the richness of 2D information in the scene, our method will choose different key parameters in the tracking and loop closure modules, enabling robust performance across diverse environments. Moreover, to alleviate the cumulative pose drift in large-scale scenes, we propose a 2D-3D Gaussian based loop closure method with BEV loop detection, pose graph optimization with voxel ICP and rendering loss. We also design an online distillation method for submap fusion that is triggered upon loop closure detection. This method merges the multiple submaps involved in the loop, thereby enhancing global consistency of the global map. Overall, our contributions are shown as follows:
  • We propose VPGS-SLAM, the first 3DGS-based large-scale RGBD SLAM framework that enables efficient and accurate scene reconstruction and pose estimation in both indoor and city-scale environments.
@@
  • A novel 2D-3D fusion tracking method is proposed to combine the 3D Gaussian geometric information with 2D photometric information for accurate pose estimation. We also propose a 2D-3D Gaussian loop closure method with BEV loop detection, and pose graph optimization. A submap fusion method with online distillation is proposed to achieve global consistency. Experiments on various indoor and outdoor datasets demonstrate the superiority of the proposed method in both mapping and tracking.
2 Related Work
@@
Traditional SLAM. SLAM has been an active research field for the past two decades. Traditional visual SLAM algorithms orbslam estimate accurate camera poses and represent the scene using sparse point clouds. DTAM dtam was the first RGB-D approach to achieve dense scene reconstruction. Some learning-based methods droidslam ; dvlo , integrate traditional geometric frameworks with deep learning networks for improved camera tracking and mapping. For 3D LiDAR odometry and mapping, similar to feature-matching methods widely used in visual SLAM, the seminal work LOAM loam proposes extracting sparse planar or edge feature points from the scan point cloud and registering them to the previous frame or the feature point map using ICP. Recently, CT-ICP cticp and KISS-ICP kissicp have achieved robust LiDAR odometry performance without the need for feature point extraction.
@@
NeRF-based SLAM. With the introduction of NeRFNeRF , iMAPimap pioneered the use of a single multi-layer perceptron (MLP) to represent the scene, while NICE-SLAM niceslam introduced learnable hierarchical feature grids. ESLAM eslam and Co-SLAM coslam further enhanced scene representation using tri-planes and joint coordinate-parametric encoding. Some methods plgslam ; incremental proposed a novel progressive scene representation that dynamically allocates new local representations. Go-SLAM goslam , Loopy-SLAM loopyslam use loop closure to enhance the camera tracking performance. pointslam ; glorieslam ; nerfloam ; pinslam use neural point-based neural radiance fields for large-scale scenes and high-accuracy reconstruction. Unlike these methods, which use neural implicit features, our approach adopts an explicit 3D Gaussian representation, significantly improving the scalability of our method.
@@
GS-based SLAM. Recently, 3D Gaussian Splatting  3dgs has emerged using 3D Gaussians as primitives for real-time neural rendering. SplaTAM splatam , MonoGS monogs , Gaussian-SLAM gaussianslam , and other works gsslam ; compactslam ; loopsplat ; gs-icp ; splatslam ; rtgslam ; sgsslam ; mg-slam ; densesplat are the pioneer works that successfully combine the advantages of 3D Gaussian Splatting with SLAM. These methods achieve fast rendering speed and high-fidelity reconstruction performance. GigaSLAM gigaslam is the concurrent work of our method which use depth estimation method for dense RGB SLAM in large scenes. However, the memory and storage usage are intensive in these GS-based SLAM systems, which makes them difficult to use in large-scale and long-sequence scenarios.
3 Method
@@
The pipeline of our system is shown in Fig. 2. The inputs of this framework are RGB frames and 3D points $\{I_{i},P_{i}\}_{i=1}^{M}$ with known camera intrinsics $K\in R_{3\times 3}$. Our model predicts camera poses $\{R_{i}|t_{i}\}^{M}_{i=1}$, color $\mathbf{c}$, and a structured 3D Gaussian scene representation. The system consists of three main modules: (i) Voxel-based progressive scene representation (Sec. 3.1), (ii) 2D-3D fusion camera tracking (Sec. 3.2), (iii) 3D Gaussian-based loop closure (Sec. 3.3). The network is incrementally updated with the system operation.
3.1 Voxel-based Progressive Scene Representation
@@
Sequentialized Multi-Resolution Voxel-based Scene Representation Although there are some voxel-based mapping methods, such as  scaffoldgs , the existing method is not well-suited to the incremental nature of SLAM systems. To better accommodate this characteristic, we further reformulated our approach into a sequentialized multi-resolution voxel-based mapping framework. When a keyframe arrives, the regions observed by this frame is voxelized with the point cloud $\boldsymbol{P}_{i}\in\mathcal{R}^{N\times 3}$, $i$ denotes the ID of the current frame. We use $\mathbf{V_{i}}\in\mathcal{R}^{N^{\prime}\times 3}$ to denote voxel centers. The center of each voxel is initialized as an anchor point $\textbf{x}_{i}^{a}\in\mathcal{R}^{3}$. Each anchor is characterized by its attributes $\mathcal{A}_{i}=\left\{\boldsymbol{f}_{i}^{a}\in\mathbb{R}^{32},\boldsymbol{l_{i}}\in\mathcal{R}^{3},\mathcal{O}_{i}\in\mathcal{R}^{k\times 3}\right\}$, where each component represents the anchor feature, scaling, and offsets, respectively. Then, we derive 3D Gaussians attributes from anchor points. The attributes of a neural Gaussian are defined as: position $\mu_{i}\in\mathcal{R}^{3}$, opacity $\alpha_{i}\in\mathcal{R}$, quaternion $q_{i}\in\mathcal{R}^{4}$, scaling $s_{i}\in\mathcal{R}^{3}$, and color $c_{i}\in\mathcal{R}^{3}$. The positions of the corresponding $k$ 3D Gaussians of current frame $I_{i}$ are calculated as:
@@
where $\{\mathcal{O}_{i}^{m}\}_{m=0}^{k-1}\in\mathcal{R}^{k\times 3}$ are the learnable offsets and $\boldsymbol{l}_{i}$ is the scaling factor associated with anchor. Then, the attributes of the Gaussians are decoded from the anchor feature $\boldsymbol{f}^{a}$, the viewing distance $\mathbf{\delta}_{i}$, $d_{i}$ through individual MLPs:
$\{\boldsymbol{f}_{i}^{a},\delta_{i},d_{i}\}\mapsto\{\{\alpha_{i}^{m}\}_{m=0}^{k-1},\{q_{i}^{m}\}_{m=0}^{k-1},\{s_{i}^{m}\}_{m=0}^{k-1},\{c_{i}^{m}\}_{m=0}^{k-1}\},\quad\delta_{i}=\left\|\mathbf{x}^{a}_{i}-\mathbf{x}^{c}_{i}\right\|_{2},\quad d_{i}=\frac{\mathbf{x}_{i}^{a}-\mathbf{x}_{i}^{c}}{\left\|\mathbf{x}_{i}^{a}-\mathbf{x}_{i}^{c}\right\|_{2}}$ (2)
where $\mathbf{x}^{c}_{i}$ denotes camera position of current frame $I_{i}$. In order to improve the efficiency, we use multiple levels voxel size based on camera distance, from fine to coarse. All attributes are decoded in a single pass. We achieve online reconstruction through the sequentialized voxel mapping method, where new voxels are continuously initialized, registered, aligned, and iteratively updated, ultimately enabling the reconstruction of the entire sequence.
@@
$\mathcal{L}_{m}=\mathcal{L}_{c}+\mathcal{L}_{\mathrm{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\mathrm{vol}}\mathcal{L}_{\mathrm{vol}}\vspace{-0.1cm}$ (4)
@@
$\vspace{-0.1cm}\mathcal{L}_{\mathrm{SSIM}}=\left(1-\lambda_{SSIM}\right)\cdot|\hat{I}-I|_{1}+\lambda_{SSIM}(1-\operatorname{SSIM}(\hat{I},I)),\quad\mathcal{L}_{\mathrm{vol}}=\sum_{i=1}^{N_{\mathrm{ng}}}\operatorname{Prod}\left(s_{i}\right)$ (5)
@@
3.2 2D-3D Fusion Camera Tracking
Most existing 3DGS methods rely on single-modality supervision, typically using rendering loss. However, our findings indicate that 2D photometric loss performs better in indoor and simulated environments, whereas 3D geometric information are more beneficial for outdoor scene reconstruction with motion blur and exposure change. To this end, we design a coarse-to-fine pose estimation, utilizing both 3D Gaussian information and 2D rendering information for both indoor and large-scale outdoor scenes. In coarse-level optimization, we use 2D photometric information $\mathcal{L}_{c},\mathcal{L}_{d}$ to optimize the camera pose, which provides a good initial estimate. In the fine-level stage, we leverage 3D geometric information to further refine the pose. However, directly applying 2D-3D optimization is not always suitable for diverse indoor and outdoor environments. To address this, we design a dynamic adaptive method to dynamically adjust the reliance on 2D/3D information during tracking. We assess the quality of rendering results with rendering loss $\mathcal{L}_{c},\mathcal{L}_{d}$ and input frames $I_{i}$. If the 2D information quality lower than the threshold $\zeta$, we will discard the coarse-level optimization and directly employ 3D Gaussian-based voxel ICP for pose refinement. This strategy significantly improves robustness in large-scale outdoor scenes, where visual inputs may suffer from illumination changes, exposure, and motion blur, enabling the system to rely more heavily on 3D geometric information when necessary. In the fine-level optimization, we incorporate a voxel-based 3D Gaussian ICP method, inspired bykissicp . We perform frame-to-map registration with the voxel-based 3D Gaussian map and adopt a double downsampling strategy, which retains only a single original point per voxel. Compared to most voxelization strategies that select the center of each voxel for downsampling, we find it more advantageous to retain the original point coordinates during the mapping process, which can avoid discretization errors. For the input points $\mathcal{S}$, we transform it into the global coordinate frame using the previous pose estimate $\{R_{i-1},t_{i-1}\}$ and predicted relative pose$\{R_{\text{pred}},t_{\text{pred}}\}$, resulting in the source points:
@@
where $\boldsymbol{P}$ denotes the scan in the local frame. We obtain a set of correspondences between the point cloud $\mathcal{S}$ and the local map =$\mathcal{M}=\left\{\boldsymbol{m}_{i}\mid\boldsymbol{m}_{i}\in\mathbb{R}^{3}\right\}$ through nearest neighbor(NN) search over the 3D Gaussian voxel map considering only correspondences with a point-to-point distance below threshold. To compute the current pose registration, we perform a robust optimization minimizing the residuals:
$\Delta\{R,t\}_{\text{est }}=\underset{\{R,t\}}{\operatorname{argmin}}\sum_{(s,q)\in\mathcal{C}\left(\tau_{t}\right)}\rho\left(\|\{R,t\}\boldsymbol{s}-\boldsymbol{m}\|_{2}\right),\quad\rho(e)=\frac{e^{2}/2}{\sigma_{t}/3+e^{2}}\vspace{-0.1cm}$ (7)
@@
NICE-SLAMniceslam 0.93 1.28 1.07 0.93 1.04 1.08 1.12 1.13 1.07
ESLAMeslam 0.72 0.68 0.54 0.55 0.55 0.60 0.70 0.65 0.65
Co-SLAMcoslam 0.96 1.02 0.95 0.83 0.87 0.94 1.01 0.93 0.95
Point-SLAMpointslam 0.52 0.38 0.30 0.50 0.44 1.26 0.77 0.57 0.59
PLGSLAMplgslam 0.64 0.65 0.49 0.51 0.52 0.54 0.65 0.57 0.57
Loopy-SLAMloopyslam 0.27 0.25 0.29 0.29 0.41 0.31 0.24 0.34 0.31
[…]
SplaTAMsplatam 0.33 0.44 0.33 0.51 0.29 0.29 0.35 0.75 0.41
MonoGSmonogs 0.35 0.24 0.32 0.38 0.22 0.26 0.16 0.82 0.34
Gaussian-SLAMgaussianslam 0.29 0.29 0.23 0.38 0.24 0.42 0.31 0.36 0.32
Photo-SLAMphotoslam 0.65 0.65 0.50 0.53 0.52 0.55 0.66 0.58 0.58
Loop-Splatloopsplat 0.28 0.22 0.17 0.22 0.15 0.49 0.20 0.30 0.26

### v2

@@ Abstract
3D Gaussian Splatting has recently shown promising results in dense visual SLAM. However, existing 3DGS-based SLAM methods are all constrained to small-room scenarios and struggle with memory explosion in large-scale urban scenes and long sequences. To this end, we propose VPGS-SLAM, a novel 3DGS-based large-scale RGBD SLAM framework for both indoor and outdoor scenarios. We design a novel voxel-based progressive 3D Gaussian mapping method with multiple submaps for compact and accurate scene representation in large-scale and long-sequence scenes. This allows us to scale up to arbitrary scenes and improves robustness (even under pose drifts). In addition, we propose a 2D-3D fusion camera tracking method to achieve robust and accurate camera tracking in both indoor and outdoor large-scale scenes. Furthermore, we design a 2D-3D Gaussian loop closure method to eliminate pose drift. We further propose a submap fusion method with online distillation to achieve global consistency in large-scale scenes when detecting a loop. Experiments on various indoor and outdoor datasets demonstrate the superiority and generalizability of the proposed framework. The code will be open-sourced on https://github.com/dtc111111/vpgs-slam.
@@
Figure 1: We present VPGS-SLAM, a novel large-scale SLAM framework with voxel-based progressive 3D Gaussian representation, 2D-3D assisted camera tracking and 3D Gaussian loop closure. Depicted in the middle, we demonstrate the large-scale globally consistent 3D Gaussian map built with our approach. At the top and bottom of the figure, we include zoomed-in views of the map with RGB and depth images rendered by our method, indicated by dashed blue and yellow boxes.
I Introduction
Visual Simultaneous localization and mapping (SLAM) [4, 6] has been a fundamental problem with wide applications such as autonomous driving [7], robotics [13], and remote sensing.
In intelligent transportation systems, accurate pose estimation constitutes a fundamental prerequisite for safe and reliable autonomous driving. Errors in vehicle pose estimation can directly compromise motion planning and control, leading to hazardous behaviors such as premature or unnecessary braking, unsafe following distances, and incorrect obstacle avoidance decisions. More critically, the consequences of pose misestimation extend beyond the affected vehicle itself. In real-world traffic environments, abnormal driving actions triggered by inaccurate pose estimates can influence surrounding vehicles through interaction and response, potentially propagating as cascading unsafe maneuvers, large-scale braking waves, and traffic flow instability. Such error amplification mechanisms may significantly elevate the risk of traffic accidents and degrade overall transportation safety, underscoring the essential role of robust and accurate pose estimation in autonomous driving systems.
@@ I Introduction
Several traditional methods[34, 39, 56, 44, 5, 64] have been introduced over the years. They use handcraft descriptors for image matching and represent scenes using sparse feature point maps. Due to the sparse nature of such point cloud, it is difficult for humans to understand how machines interact with the scene, and these methods cannot meet the demands of collision avoidance and motion planning. Attention then turns to dense scene reconstruction, exemplified by DTAM [35], Kintinuous [53], and ElasticFusion [54]. However, their performance is limited by high memory consumption, slow processing speeds.
@@
Following the introduction of Neural Radiance Fields (NeRF), numerous research efforts have focused on combining implicit scene representation with SLAM systems and autonomous driving [38]. iMAP [46] pioneered the use of a single MLP to represent the scene, while [68, 20, 51, 14] have further enhanced scene representation through hybrid feature grids, axis-aligned feature planes, joint coordinate-parametric encoding, and multiple implicit submaps. NeSLAM [16] use a depth completion and denoising method to improve scene representation. To further improve rendering speed, recent methods have started to explore 3D Gaussian Splatting (3DGS) [22] in SLAM systems, as demonstrated by [21, 31]. GS-based SLAM methods leverage a point-based representation associated with 3D Gaussian attributes and adopt the rasterization pipeline to render the images, achieving fast rendering speed and promising image quality.
@@
However, existing SLAM systems primarily focus on small-scale indoor environments; they face significant challenges in representing large-scale scenes (e.g., multi-room apartments and urban scenes). Some We outline the key challenges for indoor and outdoor large-scale 3DGS-based SLAM systems: a) Redundant 3D Gaussian ellipsoids: existing methods employ a substantial number of 3D Gaussian ellipsoids to represent the scene. These methods typically require more than 500MB to represent a small room-scale scene, which severely limits their applicability and results in memory overload in large-scale environments. b) Accumulation of errors and pose drift: Existing works rely on rendered loss for camera tracking, which proves inaccurate and unreliable in large-scale real-world environments with dramatic movement, motion blur and exposure change. c) Global inconsistency: In long sequences and large-scale scenes, when the robot revisits the same location, it is essential to ensure spatial correlations, and long-term memory for global consistency.
@@
To this end, we propose VPGS-SLAM, a novel 3DGS-based large-scale SLAM framework with voxel-based progressive 3D Gaussian representation, 2D-3D fusion camera tracking, and 2D-3D Gaussian based loop detection and correction, and submap fusion. We use a collection of submaps to represent the entire scene, which dynamically initializes local scene representation when the camera moves to the bounds of the local submaps. The entire scene is divided into multiple local submaps, which can significantly improve the scene representation capacity of large-scale scenes. The submap parameters do not need to be retained in memory after optimization, which can significantly reduce online memory requirements and enhances the scalability of the framework.
@@
For camera tracking, we propose a 2D-3D fusion tracking method to combine the 3D Gaussian ellipsoids geometric information with 2D photometric information. We optimize the pose with 2D photometric loss using RGB and depth loss in the coarse-level optimization stage. We then incorporate a 3D voxel based ICP method to perform frame-to-map pose estimation in the fine-level stage. We provide a good initial estimate for 3D point matching through the 2D rendering loss, enabling further accurate pose estimation. To handle both indoor and outdoor environments, we propose an adaptive 2D information assessing and parameter selection strategy. By assessing the richness of 2D information in the scene, our method will choose different key parameters in the tracking and loop closure modules, enabling robust performance across diverse environments. Moreover, to alleviate the cumulative pose drift in large-scale scenes, we propose a 2D-3D Gaussian based loop closure method with loop detection, pose graph optimization with voxel ICP and rendering loss. We also design an online distillation method for submap fusion that is triggered upon loop closure detection. This method merges the multiple submaps involved in the loop, thereby enhancing global consistency of the global map. Overall, our contributions are shown as follows:
  • We propose VPGS-SLAM, a novel 3DGS-based large-scale RGBD SLAM framework that enables efficient and accurate scene reconstruction and pose estimation in both indoor and city-scale environments.
@@
  • A novel 2D-3D fusion tracking method is proposed to combine the 3D Gaussian geometric information with 2D photometric information for accurate pose estimation. We also propose a 2D-3D Gaussian loop closure method with loop detection, and pose graph optimization. A submap fusion method with online distillation is proposed to achieve global consistency. Experiments on various indoor and outdoor datasets demonstrate the superiority of the proposed method in both mapping and tracking.
II Related Work
@@
Traditional SLAM. SLAM has been an active research field for the past two decades. Traditional visual SLAM algorithms [33] estimate accurate camera poses and represent the scene using sparse point clouds. [55] utilizes tightly-coupled LiDAR-Visual-Inertial odometry with multi-modal semantic information to enhance the robustness and accuracy of SLAM. DTAM [35] was the first RGB-D approach to achieve dense scene reconstruction. Some learning-based methods [47, 27], integrate traditional geometric frameworks with deep learning networks for improved camera tracking and mapping.For 3D LiDAR odometry and mapping, similar to feature-matching methods widely used in visual SLAM, the seminal work LOAM [62] proposes extracting sparse planar or edge feature points from the scan point cloud and registering them to the previous frame or the feature point map using ICP. Recently, CT-ICP [9] and KISS-ICP [49] have achieved robust LiDAR odometry performance without the need for feature point extraction.
@@
NeRF-based SLAM. With the introduction of NeRF[32], iMAP[46] pioneered the use of a single multi-layer perceptron (MLP) to represent the scene, while NICE-SLAM [68] introduced learnable hierarchical feature grids. ESLAM [20] and Co-SLAM [51] further enhanced scene representation using tri-planes and joint coordinate-parametric encoding. Some methods [14, 15, 52] proposed a novel progressive scene representation that dynamically allocates new local representations. Go-SLAM [63], Loopy-SLAM [26] use loop closure to enhance the camera tracking performance. SNI-SLAM [66] leverages semantic information. [42, 61, 10, 36] use neural point-based neural radiance fields for large-scale scenes and high-accuracy reconstruction. DDN-SLAM [23] focus on dynamic scene representation with masks. Unlike these methods, which use neural implicit features, our approach adopts an explicit 3D Gaussian representation, significantly improving the scalability of our method.
@@
GS-based SLAM. Recently, 3D Gaussian Splatting  [22] has emerged using 3D Gaussians as primitives for real-time neural rendering. SplaTAM [21], MonoGS [31], Gaussian-SLAM [60], and other works [57, 12, 65, 18, 43, 37, 25, 28, 24] are the pioneer works that successfully combine the advantages of 3D Gaussian Splatting with SLAM. These methods achieve fast rendering speed and high-fidelity reconstruction performance. GigaSLAM [11] is the concurrent work of our method which use depth estimation method for dense RGB SLAM in large scenes. However, the memory and storage usage are intensive in these GS-based SLAM systems, which makes them difficult to use in large-scale and long-sequence scenarios.
III Method
@@
The pipeline of our system is shown in Fig. 2. The inputs of this framework are RGB frames and 3D points $\{I_{i},P_{i}\}_{i=1}^{M}$ with known camera intrinsics $K\in R_{3\times 3}$. Our model predicts camera poses $\{R_{i}|t_{i}\}^{M}_{i=1}$, color $\mathbf{c}$, and a structured 3D Gaussian scene representation. The system consists of three main modules: (i) Voxel-based progressive scene representation (Sec. III-A), (ii) 2D-3D fusion camera tracking (Sec. III-B), (iii) 3D Gaussian-based loop closure (Sec. III-C). The network is incrementally updated with the system operation.
III-A Voxel-based Progressive Scene Representation
@@
Sequentialized Multi-Resolution Voxel-based Scene Representation Although there are some voxel-based mapping methods, such as  [30], the existing method is not well-suited to the incremental nature of SLAM systems. To better accommodate this characteristic, we further reformulated our approach into a sequentialized multi-resolution voxel-based mapping framework. When a keyframe arrives, the regions observed by this frame is voxelized with the point cloud $\boldsymbol{P}_{i}\in\mathcal{R}^{N\times 3}$, $i$ denotes the ID of the current frame. We use $\mathbf{V_{i}}\in\mathcal{R}^{N^{\prime}\times 3}$ to denote voxel centers. The center of each voxel is initialized as an anchor point $\textbf{x}_{i}^{a}\in\mathcal{R}^{3}$. Each anchor is characterized by its attributes $\mathcal{A}_{i}=\left\{\boldsymbol{f}_{i}^{a}\in\mathbb{R}^{32},\boldsymbol{l_{i}}\in\mathcal{R}^{3},\mathcal{O}_{i}\in\mathcal{R}^{k\times 3}\right\}$, where each component represents the anchor feature, scaling, and offsets, respectively. Then, we derive 3D Gaussians attributes from anchor points. The attributes of a neural Gaussian are defined as: position $\mu_{i}\in\mathcal{R}^{3}$, opacity $\alpha_{i}\in\mathcal{R}$, quaternion $q_{i}\in\mathcal{R}^{4}$, scaling $s_{i}\in\mathcal{R}^{3}$, and color $c_{i}\in\mathcal{R}^{3}$. The positions of the corresponding $k$ 3D Gaussians of current frame $I_{i}$ are calculated as:
@@
[…]
$\{\boldsymbol{f}_{i}^{a},\delta_{i},d_{i}\}\mapsto\{\{\alpha_{i}^{m}\}_{m=0}^{k-1},\{q_{i}^{m}\}_{m=0}^{k-1},\{s_{i}^{m}\}_{m=0}^{k-1},\{c_{i}^{m}\}_{m=0}^{k-1}\}$ (2)
$\delta_{i}=\left\|\mathbf{x}^{a}_{i}-\mathbf{x}^{c}_{i}\right\|_{2},\quad d_{i}=\frac{\mathbf{x}_{i}^{a}-\mathbf{x}_{i}^{c}}{\left\|\mathbf{x}_{i}^{a}-\mathbf{x}_{i}^{c}\right\|_{2}}$ (3)
where $\mathbf{x}^{c}_{i}$ denotes camera position of current frame $I_{i}$. The core MLPs include the opacity MLP , the color MLP and the covariance MLP. All of these F* are implemented in a LINEAR $\mapsto$ RELU $\mapsto$ LINEAR style with the hidden dimension of 32. Each branch’s output is activated with a head layer. For opacity, the output is activated by Tanh, where value 0 serves as a natural threshold for selecting valid samples and the final valid values can cover the full range of [0,1). For color, we activate the output with Sigmoid function:
$\left\{c_{0},\ldots,c_{k-1}\right\}=\operatorname{Sigmoid}\left(F_{c}\right)$ (4)
which canstrains the color into a range of (0,1). For rotation, we follow 3D-GS and activate it with a normalization to obtain a valid quaternion. For scaling, we adjust the base scaling of each anchor with the MLP output as follows:
$\left\{s_{0},\ldots,s_{k-1}\right\}=\operatorname{Sigmoid}\left(F_{s}\right)\cdot s_{v}$ (5)
Figure 3: Rendering Performance on Replica dataset [45] compared with SplaTAM [21] and Loop-splat [65].
In order to improve the efficiency, we use multiple levels voxel size based on camera distance, from fine to coarse. All attributes are decoded in a single pass. We achieve online reconstruction through the sequentialized voxel mapping method, where new voxels are continuously initialized, registered, aligned, and iteratively updated, ultimately enabling the reconstruction of the entire sequence.
@@
$\mathcal{L}_{m}=\mathcal{L}_{c}+\mathcal{L}_{\mathrm{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\mathrm{vol}}\mathcal{L}_{\mathrm{vol}}$ (7)
@@
$\mathcal{L}_{\mathrm{SSIM}}=\left(1-\lambda_{SSIM}\right)\cdot|\hat{I}-I|_{1}+\lambda_{SSIM}(1-\operatorname{SSIM}(\hat{I},I))$ (8)
$\mathcal{L}_{\mathrm{vol}}=\sum_{i=1}^{N_{\mathrm{ng}}}\operatorname{Prod}\left(s_{i}\right)$ (9)
@@
III-B 2D-3D Fusion Camera Tracking
Most existing 3DGS methods rely on single-modality supervision, typically using rendering loss. However, our findings indicate that 2D photometric loss performs better in indoor and simulated environments, whereas 3D geometric information are more beneficial for outdoor scene reconstruction with motion blur and exposure change. To this end, we design a coarse-to-fine pose estimation, utilizing both 3D Gaussian information and 2D rendering information for both indoor and large-scale outdoor scenes. In coarse-level optimization, we use 2D photometric information $\mathcal{L}_{c},\mathcal{L}_{d}$ to optimize the camera pose, which provides a good initial estimate. In the fine-level stage, we leverage 3D geometric information to further refine the pose. However, directly applying 2D-3D optimization is not always suitable for diverse indoor and outdoor environments. To address this, we design a dynamic adaptive method to dynamically adjust the reliance on 2D/3D information during tracking. We assess the quality of rendering results with rendering loss $\mathcal{L}_{c},\mathcal{L}_{d}$ and input frames $I_{i}$. If the 2D information quality lower than the threshold $\zeta$, we will discard the coarse-level optimization and directly employ 3D Gaussian-based voxel ICP for pose refinement. This strategy significantly improves robustness in large-scale outdoor scenes, where visual inputs may suffer from illumination changes, exposure, and motion blur, enabling the system to rely more heavily on 3D geometric information when necessary.
[…]
@@
where $\boldsymbol{P}$ denotes the scan in the local frame. We obtain a set of correspondences between the point cloud $\mathcal{S}$ and the local map $\mathcal{M}=\left\{\boldsymbol{m}_{i}\mid\boldsymbol{m}_{i}\in\mathbb{R}^{3}\right\}$ through nearest neighbor(NN) search over the 3D Gaussian voxel map considering only correspondences with a point-to-point distance below threshold. To compute the current pose registration, we perform a robust optimization minimizing the residuals:
$\Delta\{R,t\}_{\text{est }}=\underset{\{R,t\}}{\operatorname{argmin}}\sum_{(s,q)\in\mathcal{C}\left(\tau_{t}\right)}\rho\left(\|\{R,t\}\boldsymbol{s}-\boldsymbol{m}\|_{2}\right)$ (11)
$\rho(e)=\frac{e^{2}/2}{\sigma_{t}/3+e^{2}}$ (12)
@@
NICE-SLAM[68] 0.93 1.28 1.07 0.93 1.04 1.08 1.12 1.13 1.07
ESLAM[20] 0.72 0.68 0.54 0.55 0.55 0.60 0.70 0.65 0.65
Co-SLAM[51] 0.96 1.02 0.95 0.83 0.87 0.94 1.01 0.93 0.95
Point-SLAM[42] 0.52 0.38 0.30 0.50 0.44 1.26 0.77 0.57 0.59
PLGSLAM[14] 0.64 0.65 0.49 0.51 0.52 0.54 0.65 0.57 0.57
Loopy-SLAM[26] 0.27 0.25 0.29 0.29 0.41 0.31 0.24 0.34 0.31
3DGS-based Methods
SplaTAM[21] 0.33 0.44 0.33 0.51 0.29 0.29 0.35 0.75 0.41
MonoGS[31] 0.35 0.24 0.32 0.38 0.22 0.26 0.16 0.82 0.34
Gaussian-SLAM[60] 0.29 0.29 0.23 0.38 0.24 0.42 0.31 0.36 0.32
Photo-SLAM[19] 0.65 0.65 0.50 0.53 0.52 0.55 0.66 0.58 0.58
Loop-Splat[65] 0.28 0.22 0.17 0.22 0.15 0.49 0.20 0.30 0.26